# Computation of Robust Option Prices via Structured Multi-Marginal Martingale Optimal Transport

### v1

@@ 4.2 Exploiting the structure for computing the projections
The computational bottleneck of Algorithm 1 is the evaluation of the projections; this is because the number of elements in the tensors grows exponentially in the number of marginals. A key ingredient for efficiently computing the optimal dual variables is therefore to reduce the work required for this part. We will now show how this can be done by exploiting the structure inherent in the problem, hence simplifying equations (30) and (31).
@@
Let $\lambda=\{\lambda_{t}\}_{t\in\mathcal{T}}$ and $\gamma=\{\gamma_{t}\}_{t=0}^{T-1}$ be given families of vectors in $\mathbb{R}^{n_{S}}$ and $\mathbb{R}^{n_{S}n_{X}}$, respectively. Let $u_{t}=\exp(\lambda_{t}/\epsilon)$ for $t\in{\mathcal{T}}$ and let $K_{t}=\exp(-C_{t}/\epsilon)$ and $G_{t}=\exp(\operatorname{diag}(\gamma_{t})\Delta/\epsilon)$ for $t\in[T]\backslash 0$, where the matrices $C_{t}$, $t\in[T]\backslash 0$, are as in problem (22) and the matrix $\Delta$ is as given in equation 19. Define two families of vectors, $\hat{\psi}=\{\hat{\psi}_{t}\}_{t=0}^{T}$ and $\psi=\{\psi_{t}\}_{t=0}^{T}$, via the recursions
$\hat{\psi}_{t}=\begin{cases}\mathbf{1}_{n_{S}n_{X}},&t=0\\ \left(K_{t}\odot G_{t}\right)^{\top}(\hat{\psi}_{t-1}\odot(\mathbf{1}_{n_{X}}\otimes u_{t-1})),&(t-1)\in\mathcal{T},\quad t\in[T]\backslash 0\\ \left(K_{t}\odot G_{t}\right)^{\top}\hat{\psi}_{t-1},&(t-1)\not\in\mathcal{T},\quad t\in[T]\backslash 0\end{cases}$ (32)
@@
$\psi_{t}=\begin{cases}\mathbf{1}_{n_{S}n_{X}},&t=T\\ \left(K_{t+1}\odot G_{t+1}\right)\left(\psi_{t+1}\odot(\mathbf{1}_{n_{X}}\otimes u_{t+1})\right),&(t+1)\in\mathcal{T},\quad t\in[T-1]\\ \left(K_{t+1}\odot G_{t+1}\right)\psi_{t+1},&(t+1)\not\in\mathcal{T},\quad t\in[T-1].\end{cases}$ (33)
Then $\lambda$ and $\gamma$ are optimal variables for the dual problem (28) if and only if the following equations hold
$u_{t}=\upmu_{t}\oslash P^{S}(\hat{\psi}_{t}\odot\psi_{t}),\quad t\in\mathcal{T},$
@@
$\hat{\psi}_{t}\odot(\mathbf{1}_{n_{X}}\otimes u_{t})\odot\left(K_{t+1}\odot G_{t+1}\odot\Delta\right)\left(\psi_{t+1}\odot(\mathbf{1}_{n_{X}}\otimes u_{t+1})\right)=\mathbf{0}_{n_{S}n_{X}},$ $t\in\mathcal{T},\;$ $t+1\in\mathcal{T}$
$\hat{\psi}_{t}\odot\left(K_{t+1}\odot G_{t+1}\odot\Delta\right)\left(\psi_{t+1}\odot(\mathbf{1}_{n_{X}}\otimes u_{t+1})\right)=\mathbf{0}_{n_{S}n_{X}},$ $t\notin\mathcal{T},\;$ $t+1\in\mathcal{T}$
$\hat{\psi}_{t}\odot(\mathbf{1}_{n_{X}}\otimes u_{t})\odot\left(K_{t+1}\odot G_{t+1}\odot\Delta\right)\psi_{t+1}=\mathbf{0}_{n_{S}n_{X}},$ $t\in\mathcal{T},\;$ $t+1\notin\mathcal{T}$
$\hat{\psi}_{t}\odot\left(K_{t+1}\odot G_{t+1}\odot\Delta\right)\psi_{t+1}=\mathbf{0}_{n_{S}n_{X}},$ $t\notin\mathcal{T},\;$ $t+1\notin\mathcal{T}.$
@@
According to the above theorem, we can compute dual variables satisfying the optimality conditions from equations (29a) and (29b), by simply performing a number of matrix-vector products. This greatly reduces the computational work required, especially since the help vectors $\hat{\psi}$ (resp. $\psi$) can be computed inductively (resp. recursively); therefore, we do not have to recalculate the full chains for each individual variable update. We also emphasise that according to the above theorem, there is no need to explicitly form the tensors $K$, $U_{\lambda}$ and $G_{\gamma}$.
In order to prove Theorem 4.1, we will need the following lemma.
@@
Let $\lambda=\{\lambda_{t}\}_{t\in\mathcal{T}}$ and $\gamma=\{\gamma_{t}\}_{t=0}^{T-1}$ be given families of vectors in $\mathbb{R}^{n_{S}}$ and $\mathbb{R}^{n_{S}n_{X}}$, respectively. Define $u_{t}$ for $t\in\mathcal{T}$ and $K_{t}$ and $G_{t}$ for $t\in[T]\backslash 0$ as in Theorem 4.1 and let $U_{\lambda}(\boldsymbol{\omega})=\sum_{t\in\mathcal{T}}u_{t}(\omega_{t}^{S})$ and $G_{\gamma}(\boldsymbol{\omega})=\sum_{t\in[T]\backslash 0}G_{t}(\omega_{t-1},\omega_{t})$ for $\boldsymbol{\omega}\in\boldsymbol{\Omega}$. Then
$P_{t}$ $\left(K\odot U_{\lambda}\odot G_{\gamma}\right)=\hat{\psi}_{t}\odot\bar{u}_{t}\odot\psi_{t},$ $t\in[T],$
where $\hat{\psi}$ and $\psi$ are functions of $\lambda$ and $\gamma$ as given in equations (32) and (33), respectively, and $\bar{u}_{t}\in\mathbb{R}^{n_{S}n_{X}}$ is given by
$\bar{u}_{t}=\begin{cases}\mathbf{1}_{n_{X}}\otimes u_{t},&t\in\mathcal{T}\\ \mathbf{1}_{n_{S}n_{X}},&t\in[T]\backslash\mathcal{T}.\end{cases}$ (34)
@@
$P_{t_{1},t_{2}}$ $\left(K\odot U_{\lambda}\odot G_{\gamma}\right)$
$=\operatorname{diag}(\hat{\psi}_{t_{1}}\odot\bar{u}_{t_{1}})\big{(}(K_{t_{1}+1}\odot G_{t_{1}+1})\operatorname{diag}(\bar{u}_{t_{1}+1})\dots(K_{t_{2}}\odot G_{t_{2}})\operatorname{diag}(\bar{u}_{t_{2}})\big{)}\operatorname{diag}(\psi_{t_{2}}).$
Proof.
Start by noting that $U_{\lambda}$ is of the form
$U_{\lambda}(\omega_{0},\dots,\omega_{T})=\prod_{t\in[T]}\bar{u}_{t}(\omega_{t}),\quad\omega_{0},\dots,\omega_{T}\in\{1,\dots,n_{S}\}\times\{1,\dots,n_{X}\}$
since, when $\omega_{t}\in\{1,\dots,n_{S}\}\times\{1,\dots,n_{X}\}$ is ordered according to (18), $\bar{u}_{t}(\omega_{t})=u_{t}(\omega_{t}^{S})$ for $t\in\mathcal{T}$. Also note that $K\odot G_{\gamma}$ inherits the structure of $K$ and $G_{\gamma}$; indeed,
$\left(K\odot G_{\gamma}\right)(\omega_{0},\dots,\omega_{T})=\prod_{t\in[T]\backslash 0}\left(K_{t}\odot G_{t}\right)(\omega_{t-1},\omega_{t}),\quad\omega_{0},\dots,\omega_{T}\in\{1,\dots,n_{S}\}\times\{1,\dots,n_{X}\}.$
The requirements of Proposition 2.8 are thus satisfied and it yields expressions for the projections. Specifically, taking
$\hat{\psi}_{t}:=\begin{cases}\mathbf{1}_{n_{S}n_{X}},&t=0\\ \left(K_{t}\odot G_{t}\right)^{\top}\operatorname{diag}(\bar{u}_{t-1})\dots\left(K_{1}\odot G_{1}\right)^{\top}\bar{u}_{0},&t\in[T]\backslash 0\end{cases}$
and
$\psi_{t}:=\begin{cases}\mathbf{1}_{n_{S}n_{X}},&t=T\\ \left(K_{t+1}\odot G_{t+1}\right)\operatorname{diag}(\bar{u}_{t+1})\dots\left(K_{T}\odot G_{T}\right)\bar{u}_{T},&t\in[T-1]\end{cases}$
we get
$P_{t}\left(K\odot U_{\lambda}\odot G_{\gamma}\right)=\hat{\psi}_{t}\odot\bar{u}_{t}\odot\psi_{t},\quad t\in[T].$
It is easily verified that this choice of $\hat{\psi}$ and $\psi$ are given in equations (32) and (33), respectively, given the definition from equation 34.
Let $t_{1},t_{2}\in[T]$ such that $t_{1}<t_{2}$. Applying the results from [23, Lemma 2 and proof of Proposition 2] and making the same arguments as for the one-marginal projection gives that
$P_{t_{1},t_{2}}(K\odot U_{\lambda}\odot G_{\gamma})=\operatorname{diag}(\hat{\psi}_{t_{1}}\odot\bar{u}_{t_{1}})M_{t_{1},t_{2}}\operatorname{diag}(\bar{u}_{t_{2}}\odot\psi_{t_{2}})$
where
$M_{t_{1},t_{2}}=(K_{t_{1}+1}\odot G_{t_{1}+1})\operatorname{diag}(\bar{u}_{t_{1}+1})\dots(K_{t_{2}-1}\odot G_{t_{2}-1})\operatorname{diag}(\bar{u}_{t_{2}-1})(K_{t_{2}}\odot G_{t_{2}}),$
from where the result follows. ∎
@@
It was shown in Section 4.1 that the (representation of) the optimal dual variables $u$ and $\gamma$ should satisfy the optimality conditions in equations (30) and (31). Start by considering the former equation, whose right-hand side is equal to
$\left(\upmu_{t}\odot u_{t}\right)\oslash P_{t}^{S}(K\odot U_{\lambda}\odot G_{\gamma}),\quad t\in\mathcal{T}.$ (35)
We must find an expression for the projection of the tensor $K\odot U_{\lambda}\odot G_{\gamma}$. Fix some $t\in\mathcal{T}$ and start by recalling that $P_{t}^{S}(K\odot U_{\lambda}\odot G_{\gamma})=(P^{S}\circ P_{t})(K\odot U_{\lambda}\odot G_{\gamma})$, where Lemma 4.3 then yields an expression for $P_{t}(K\odot U_{\lambda}\odot G_{\gamma})$. Therefore
$P_{t}^{S}\left(K\odot U_{\lambda}\odot G_{\gamma}\right)=P^{S}\big{(}\hat{\psi}_{t}\odot\left(\mathbf{1}_{n_{X}}\otimes u_{t}\right)\odot\psi_{t}\big{)}=P^{S}\big{(}\hat{\psi}_{t}\odot\psi_{t}\big{)}\odot u_{t}$ (36)
where the second equality follows from the fact that $\left(\mathbf{1}_{n_{X}}\otimes u_{t}\right)(\omega_{t})=u_{t}(\omega_{t}^{S})$ for every $\omega_{t}\in\{1,\dots,n_{S}\}\times\{1,\dots,n_{X}\}$ ordered by (18). Inserting equation 36 into equation 35 yields the expression stated in the theorem.
Moving on to equation 31, we apply Lemma 4.3 to obtain an expression for the bi-marginal projections,
$P_{t,t+1}\left(K\odot U_{\lambda}\odot G_{\gamma}\right)=\operatorname{diag}(\hat{\psi}_{t}\odot\bar{u}_{t})(K_{t+1}\odot G_{t+1})\operatorname{diag}(\psi_{t+1}\odot\bar{u}_{t+1}),\quad t\in[T-1].$
The optimality condition from equation 31 thus becomes
$\hat{\psi}_{t}\odot\bar{u}_{t}\odot\left(K_{t+1}\odot G_{t+1}\odot\Delta\right)(\psi_{t+1}\odot\bar{u}_{t+1})$ $=\mathbf{0}_{n_{S}n_{X}},\quad t\in[T-1].$
@@
We end this section by noting that Lemma 4.3 provides an expression that allows for fast computation of the sub-transport between any two marginals, without explicitly having to form the full transport $Q$. One type of sub-transport that is of particular interest is transportation between adjacent marginals; later we will see that such sub-transports allow for recovering the price without explicitly forming the full transport $Q$.
@@
Let $\lambda=\{\lambda_{t}\}_{t\in\mathcal{T}}$ and $\gamma=\{\gamma_{t}\}_{t=0}^{T-1}$ be given families of vectors in $\mathbb{R}^{n_{S}}$ and $\mathbb{R}^{n_{S}n_{X}}$, respectively. Define $u_{t}$ for $t\in\mathcal{T}$ and $K_{t}$ and $G_{t}$ for $t\in[T]\backslash 0$ as in Theorem 4.1 and let the corresponding transport plan be denoted $Q_{\lambda,\gamma}$. Then the bi-marginal sub-transport $P_{t-1,t}(Q_{\lambda,\gamma)}$ for $t\in[T]\backslash 0$ is given by
@@
Fix some $t\in[T]\backslash 0$ and note that by equation 27, $P_{t-1,t}(Q_{\lambda,\gamma})=P_{t-1,t}(K\odot U_{\lambda}\odot G_{\gamma})$, where
$P_{t-1,t}(K\odot U_{\lambda}\odot G_{\gamma})=\operatorname{diag}\big{(}\hat{\psi}_{t-1}\odot\bar{u}_{t-1}\big{)}\left(K_{t}\odot G_{t}\right)\operatorname{diag}\big{(}\psi_{t}\odot\bar{u}_{t}\big{)}$
follows from application of Lemma 4.3 with $t_{1}=t-1$ and $t_{2}=t$. Here $\bar{u}_{t-1}$ and $\bar{u}_{t}$ are defined as in equation 34. This proves the assertion. ∎

### v2

@@ 4.2 Exploiting the structure for computing the projections
The computational bottleneck of Algorithm 1 is the evaluation of the projections; this is because the number of elements in the tensors grows exponentially in the number of marginals. A key ingredient for efficiently computing the optimal dual variables is therefore to reduce the work required for this part. We will now show how this can be done by exploiting the structure inherent in the problem, hence simplifying equations (25a) and (25b).
@@
Let $\lambda=\{\lambda_{t}\}_{t\in\mathcal{T}}$ and $\gamma=\{\gamma_{t}\}_{t\in[T-1]}$ be given families of vectors with $\lambda_{t}\in\mathbb{R}^{n_{t}^{\mathcal{S}}}$ and $\gamma_{t}\in\mathbb{R}^{n_{t}}$, respectively. Let $u_{t}^{\lambda_{t}}=\exp(\lambda_{t}/\varepsilon)$ for $t\in{\mathcal{T}}$ and let $K_{t}=\exp(-C_{t}/\varepsilon)$ and $G^{\gamma_{t-1}}_{t}=\exp(\operatorname{diag}(\gamma_{t-1})\Delta_{t}/\varepsilon)$, where the matrices $C_{t}$ are as in problem (20) and the matrices $\Delta_{t}$ are as given in equation 18, for $t\in[T]\backslash 0$. Define two families of vectors, $\hat{\psi}=\{\hat{\psi}_{t}\}_{t\in[T]}$ and $\psi=\{\psi_{t}\}_{t\in[T]}$, via the recursions
$\hat{\psi}_{t}=\begin{cases}1_{n_{0}},&t=0\\ (K_{t}\odot G^{\gamma_{t-1}}_{t})^{\top}(\hat{\psi}_{t-1}\odot(1_{n_{t-1}^{\mathcal{X}}}\otimes u_{t-1}^{\lambda_{t-1}})),&(t-1)\in\mathcal{T},\quad t\in[T]\backslash 0\\ (K_{t}\odot G^{\gamma_{t-1}}_{t})^{\top}\hat{\psi}_{t-1},&(t-1)\not\in\mathcal{T},\quad t\in[T]\backslash 0\end{cases}$ (28)
@@
$\psi_{t}=\begin{cases}1_{n_{T}},&t=T\\ (K_{t+1}\odot G^{\gamma_{t}}_{t+1})(\psi_{t+1}\odot(1_{n_{t+1}^{\mathcal{X}}}\otimes u_{t+1}^{\lambda_{t+1}})),&(t+1)\in\mathcal{T},\quad t\in[T-1]\\ (K_{t+1}\odot G^{\gamma_{t}}_{t+1})\psi_{t+1},&(t+1)\not\in\mathcal{T},\quad t\in[T-1].\end{cases}$ (29)
Suppose that the grid ${}_{t\in[T]}\mathcal{S}_{t}$ and the given marginals $\{\mu_{t}\}_{t\in\mathcal{T}}$ satisfy the assumptions of the second part of Theorem 4.2. Then $(\lambda,\gamma)$ are optimal variables for the dual problem (24) if and only if the following equations hold
$u_{t}^{\lambda_{t}}=m_{t}\oslash P^{\mathcal{S}}(\hat{\psi}_{t}\odot\psi_{t}),\quad t\in\mathcal{T},$
@@
$\hat{\psi}_{t}\odot\big{(}1_{n_{t}^{\mathcal{X}}}\otimes u_{t}^{\lambda_{t}}\big{)}\odot\big{(}K_{t+1}\odot G^{\gamma_{t}}_{t+1}\odot\Delta_{t+1}\big{)}\big{(}\psi_{t+1}\odot(1_{n_{t+1}^{\mathcal{X}}}\otimes u_{t+1}^{\lambda_{t+1}})\big{)}=0_{n_{t}},$ $t\in\mathcal{T},\;$ $t+1\in\mathcal{T}$
$\hat{\psi}_{t}\odot\big{(}K_{t+1}\odot G^{\gamma_{t}}_{t+1}\odot\Delta_{t+1}\big{)}\big{(}\psi_{t+1}\odot(1_{n_{t+1}^{\mathcal{X}}}\otimes u_{t+1}^{\lambda_{t+1}})\big{)}=0_{n_{t}},$ $t\notin\mathcal{T},\;$ $t+1\in\mathcal{T}$
$\hat{\psi}_{t}\odot(1_{n_{t}^{\mathcal{X}}}\otimes u_{t}^{\lambda_{t}})\odot\left(K_{t+1}\odot G^{\gamma_{t}}_{t+1}\odot\Delta_{t+1}\right)\psi_{t+1}=0_{n_{t}},$ $t\in\mathcal{T},\;$ $t+1\notin\mathcal{T}$
$\hat{\psi}_{t}\odot\left(K_{t+1}\odot G^{\gamma_{t}}_{t+1}\odot\Delta_{t+1}\right)\psi_{t+1}=0_{n_{t}},$ $t\notin\mathcal{T},\;$ $t+1\notin\mathcal{T}.$
@@
According to the above theorem, we can compute dual variables satisfying the optimality conditions from equations (25a) and (25b), by simply performing a number of matrix-vector products. This greatly reduces the computational work required, especially since the help vectors $\hat{\psi}$ (resp. $\psi$) can be computed inductively (resp. recursively); therefore, we do not have to recalculate the full chains for each individual variable update. We also emphasise that according to the above theorem, there is no need to explicitly form the tensors $\mathbf{K}$, $\mathbf{U}^{\lambda}$ and $\mathbf{G}^{\gamma}$.
In order to prove Theorem 4.5, we will need the following lemma. Its proof can be found in Appendix: Additional proofs.
@@
Let $\lambda=\{\lambda_{t}\}_{t\in\mathcal{T}}$ and $\gamma=\{\gamma_{t}\}_{t\in[T-1]}$ be given families of vectors with $\lambda_{t}\in\mathbb{R}^{n_{t}^{\mathcal{S}}}$ and $\gamma_{t}\in\mathbb{R}^{n_{t}}$, respectively. Define $u_{t}^{\lambda_{t}}$ for $t\in\mathcal{T}$ and $K_{t}$ and $G^{\gamma_{t-1}}_{t}$ for $t\in[T]\backslash 0$ as in Theorem 4.5 and let $\mathbf{K}(i_{0},\dots,i_{T})=\sum_{t\in[T]\backslash 0}K_{t}(i_{t-1},i_{t})$, $\mathbf{U}^{\lambda}(i_{0},\dots,i_{T})=\sum_{t\in\mathcal{T}}u_{t}^{\lambda_{t}}(i_{t}^{\mathcal{S}})$ and $\mathbf{G}^{\gamma}(i_{0},\dots,i_{T})=\sum_{t\in[T]\backslash 0}G^{\gamma_{t-1}}_{t}(i_{t-1},i_{t})$ for $(i_{0},\dots,i_{T})\in\mathcal{I}$. Then
$P_{t}$ $\left(\mathbf{K}\odot\mathbf{U}^{\lambda}\odot\mathbf{G}^{\gamma}\right)=\hat{\psi}_{t}\odot\bar{u}_{t}\odot\psi_{t},$ $t\in[T],$
where $\hat{\psi}$ and $\psi$ are functions of $\lambda$ and $\gamma$ as given in equations (28) and (29), respectively, and $\bar{u}_{t}\in\mathbb{R}^{n_{t}}$ is given by
$\bar{u}_{t}=\begin{cases}1_{n_{t}^{\mathcal{X}}}\otimes u_{t}^{\lambda_{t}},&t\in\mathcal{T}\\ 1_{n_{t}},&t\in[T]\backslash\mathcal{T}.\end{cases}$ (30)
@@
$P_{t_{1},t_{2}}$ $(\mathbf{K}\odot\mathbf{U}^{\lambda}\odot\mathbf{G}^{\gamma})$
$=\operatorname{diag}(\hat{\psi}_{t_{1}}\odot\bar{u}_{t_{1}})\big{(}(K_{t_{1}+1}\odot G^{\gamma_{t_{1}}}_{t_{1}+1})\operatorname{diag}(\bar{u}_{t_{1}+1})\dots(K_{t_{2}}\odot G^{\gamma_{t_{2}-1}}_{t_{2}})\operatorname{diag}(\bar{u}_{t_{2}})\big{)}\operatorname{diag}(\psi_{t_{2}}).$
@@
Theorem 4.2 states that the (representation of) the optimal dual variables $u^{\lambda}$ and $\gamma$ should satisfy the optimality conditions (25a) and (25b). Start by considering the former, whose right-hand side is equal to
$(m_{t}\odot u_{t}^{\lambda_{t}})\oslash P_{t}^{\mathcal{S}}(\mathbf{K}\odot\mathbf{U}^{\lambda}\odot\mathbf{G}^{\gamma}),\quad t\in\mathcal{T}.$ (31)
We must find an expression for the projection of the tensor $\mathbf{K}\odot\mathbf{U}^{\lambda}\odot\mathbf{G}^{\gamma}$. Fix some $t\in\mathcal{T}$ and start by recalling that $P_{t}^{\mathcal{S}}(\mathbf{K}\odot\mathbf{U}^{\lambda}\odot\mathbf{G}^{\gamma})=(P^{\mathcal{S}_{t}}\circ P_{t})(\mathbf{K}\odot\mathbf{U}^{\lambda}\odot\mathbf{G}^{\gamma})$, where Lemma 4.7 then yields an expression for $P_{t}(\mathbf{K}\odot\mathbf{U}^{\lambda}\odot\mathbf{G}^{\gamma})$. Therefore
$P_{t}^{\mathcal{S}}(\mathbf{K}\odot\mathbf{U}^{\lambda}\odot\mathbf{G}^{\gamma})=P^{\mathcal{S}_{t}}\big{(}\hat{\psi}_{t}\odot(1_{n_{t}^{\mathcal{X}}}\otimes u_{t}^{\lambda_{t}})\odot\psi_{t}\big{)}=P^{\mathcal{S}_{t}}(\hat{\psi}_{t}\odot\psi_{t})\odot u_{t}^{\lambda_{t}}$ (32)
where the second equality follows from the fact that $(1_{n_{t}^{\mathcal{X}}}\otimes u_{t}^{\lambda_{t}})(i_{t})=u_{t}^{\lambda_{t}}(i_{t}^{\mathcal{S}})$ for every $i_{t}\in\mathcal{I}_{t}$ given by the order (17). Inserting equation 32 into equation 31 yields the expression stated in the theorem.
Moving on to equation 25b, we apply Lemma 4.7 to obtain an expression for the bi-marginal projections, $P_{t,t+1}(\mathbf{K}\odot\mathbf{U}^{\lambda}\odot\mathbf{G}^{\gamma})=\operatorname{diag}(\hat{\psi}_{t}\odot\bar{u}_{t})(K_{t+1}\odot G^{\gamma_{t}}_{t+1})\operatorname{diag}(\psi_{t+1}\odot\bar{u}_{t+1})$ for $t\in[T-1]$. The optimality condition from equation 25b thus becomes
$\hat{\psi}_{t}\odot\bar{u}_{t}\odot(K_{t+1}\odot G^{\gamma_{t}}_{t+1}\odot\Delta_{t+1})(\psi_{t+1}\odot\bar{u}_{t+1})$ $=0_{n_{t}},\quad t\in[T-1].$
@@
We end this section by noting that Lemma 4.7 provides an expression that allows for fast computation of the sub-transport between any two marginals, without explicitly having to form the full transport $\mathbf{Q}$. One type of sub-transport that is of particular interest is transportation between adjacent marginals; later we will see that such sub-transports allow for recovering the price without explicitly forming the full transport $\mathbf{Q}$.
@@
Let $\lambda=\{\lambda_{t}\}_{t\in\mathcal{T}}$ and $\gamma=\{\gamma_{t}\}_{t\in[T-1]}$ be given families of vectors with $\lambda_{t}\in\mathbb{R}^{n_{t}^{\mathcal{S}}}$ and $\gamma_{t}\in\mathbb{R}^{n_{t}}$, respectively. Define $u_{t}^{\lambda_{t}}$ for $t\in\mathcal{T}$ and $K_{t}$ and $G^{\gamma_{t-1}}_{t}$ for $t\in[T]\backslash 0$ as in Theorem 4.5 and let a corresponding transport plan $\mathbf{Q}^{\lambda,\gamma}$ be given by equation 26. Then the bi-marginal sub-transport $P_{t-1,t}(\mathbf{Q}^{\lambda,\gamma})$ for $t\in[T]\backslash 0$ is given by
@@
Fix some $t\in[T]\backslash 0$ and note that by equation 26, $P_{t-1,t}(\mathbf{Q}^{\lambda,\gamma})=P_{t-1,t}(\mathbf{K}\odot\mathbf{U}^{\lambda}\odot\mathbf{G}^{\gamma})$, where $P_{t-1,t}(\mathbf{K}\odot\mathbf{U}^{\lambda}\odot\mathbf{G}^{\gamma})=\operatorname{diag}(\hat{\psi}_{t-1}\odot\bar{u}_{t-1})(K_{t}\odot G^{\gamma_{t-1}}_{t})\operatorname{diag}(\psi_{t}\odot\bar{u}_{t}^{\lambda_{t}})$ follows from application of Lemma 4.7 with $t_{1}=t-1$ and $t_{2}=t$. Here $\bar{u}_{t-1}$ and $\bar{u}_{t}$ are defined as in equation 30. This proves the assertion. ∎